Brute force attack

What is the maximum amount of time that it would take a computer that can test 1000 billion (1012) keys per second to carry out a brute-force attack on a file encrypted using 128-bit AES? Give your answer in years, using scientific notation, correct to 2 significant figures. 
Answer #1

No of possible keys for 128-bit AES =  \huge 2^{128}

This is because each bit can have 2 values (0 or 1) and therefore, number of different 128-bit keys = \huge 2 \times 2 \times2 \times .....\text{(total 128 times)} = 2^{128} . This is our search space and these many keys we would have to examine.

Our computer can test \huge 10^{12} keys per second. In other words,

Time taken to test  \huge 10^{12} keys = 1 second

\huge \text{Time taken to test 1 key }=\frac{1}{10^{12}} \text{ second}

\huge \text{Time taken to test }2^{128} \text{ keys }=\frac{2^{128}}{10^{12}} \text{ second}

As 1 year is approximately equal to \huge 3.2 \times 10^7 second (given in the question), we have,

\huge \text{Time taken to test }2^{128} \text{ keys }=\frac{2^{128}}{10^{12}\times 3.2 \times 10^7} \text{ years}

\huge =\frac{2^{128}}{10^{18}\times 32} \text{ years}

  \huge =\frac{2^{128}}{10^{18}\times 2^5} \text{ years}

  \huge =\frac{2^{123}}{10^{18}} \text{ years}

\huge =1.063 \times 10^{19} \text{ years}

\huge =\mathbf{1.1 \times 10^{19}} \text{ years, correct to two significant figures}.

So. the machine will take \huge 1.1 \times 10^{19} years to crack a 128-bit AES.